OpenStudy (accidentalaichan):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 8 -x - 3y = -12 3y + 12 -3y + 12 3y - 12 -3y - 12

OpenStudy (shamil98):

solve for x then. -x -3y = -12

OpenStudy (mertsj):

1. add 3y to both sides 2. Multiply both sides by -1

OpenStudy (accidentalaichan):

x = 4?

OpenStudy (accidentalaichan):

I'm having to find an equation

OpenStudy (shamil98):

not the value of x.

OpenStudy (shamil98):

-x - 3y = -12 -x = 3y - 12 one more step.. divide both sides by -1.

OpenStudy (accidentalaichan):

3y - 12?

OpenStudy (shamil98):

Divide everything in -x = 3y - 12 by -1.

OpenStudy (mertsj):

You're supposed to have an equation. What happened to the x and the equal sign?

OpenStudy (accidentalaichan):

None of the choices have an x and equal sign

OpenStudy (accidentalaichan):

-3y - 12

OpenStudy (shamil98):

-x = 3y - 12 x = -3y + 12

OpenStudy (shamil98):

you were close but remember -12/-1 = 12

OpenStudy (accidentalaichan):

Thanks ;-; I hate systems of equations. You're so kind